OpenStudy (anonymous):

Find all polar coordinates of point P where P =(5,pi/3)

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (anonymous):

(5, pi divided by 3 + nπ) or (-5, pi divided by 3 + nπ) (5, pi divided by 3 + (2n + 1)π) or (-5, pi divided by 3 + 2nπ) (5, pi divided by 3 + 2nπ) or (-5, pi divided by 3 + (2n + 1)π) (5, pi divided by 3 + 2nπ) or (-5, pi divided by 3 + 2nπ)

OpenStudy (ipwnbunnies):

Ok. It looks kinda confusing. Do you understand that 2pi radians makes a complete circle. And pi radians a half a circle?

OpenStudy (anonymous):

yep i understand that

OpenStudy (ipwnbunnies):

Ok. So if we have point (5, pi/3) If we add any multiple of 2pi to the pi/3, it'll just go back to the same angle, right?

OpenStudy (anonymous):

yah makes sense

OpenStudy (ipwnbunnies):

Good, so for the first part, we can write it like this: \[(5, \frac{\pi}{3} + 2 \pi n)\] Where 'n' can be any integer.

OpenStudy (anonymous):

Alright, ya that eliminates 2 options :D

OpenStudy (ipwnbunnies):

Yes! Now, for the next part.

OpenStudy (ipwnbunnies):

(5, pi/3) is the same as (-5, pi/3 + pi)

OpenStudy (ipwnbunnies):

It's the angle exactly opposite of pi/3. But since 'r' is negative, we trace it backwards, so it's the same point.

OpenStudy (anonymous):

hmmm alright. ya i understand that

OpenStudy (ipwnbunnies):

Ok, good. However, it can't be any multiple of pi. Because if we use pi/3 + 2pi, that'll be an angle similar to pi/3, which isn't what we want. We want every ODD integer, which we can write like this: \[(5, \frac{\pi}{3}) = (-5,\frac{\pi}{3} + (2n + 1)\pi)\] Where 'n' is any integer.

OpenStudy (anonymous):

Thanks man! Do you have time for one more problem? It's my last one. I really appreciate your help

OpenStudy (ipwnbunnies):

Ok.